OpenStudy (anonymous):

please help me solve the improper integral of the integral from 0 to 2 of (2x+1)/square root of(x2-x

OpenStudy (anonymous):

substitute u = x^2-x du = 2x-1

OpenStudy (anonymous):

is the denominator x^2+x or x^2-x ?

OpenStudy (anonymous):

the sign of the denominator is a plus

OpenStudy (anonymous):

then do what I suggested. u = x^2+x du = 2x+1

OpenStudy (anonymous):

so it becomes the integral from 0 to2 of du/square root of u which gives2(x2-1)1/2 from c to2?

OpenStudy (anonymous):

if u = x^2+x, then at x =0, u =0 and x =2, u = 6 so your integral becomes \[\int\limits_{0}^{6}\frac{du}{\sqrt u}\]

OpenStudy (anonymous):

thanks pls, what of reduced formulas?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

reduced formulas? what do you mean?

OpenStudy (anonymous):

fine am saying i have a problem in finding reduced formulas of trigonometric fxns in integration.

OpenStudy (anonymous):

then you should post it. vague questions will get you vague answers

OpenStudy (anonymous):

what is the reduced formula for costx power n dx?

OpenStudy (anonymous):

use equation button below to type out your question | | | v